OpenStudy (samsan9):

For the polynomial f(x)=x^3-x^2-4x-4 How many zeros does the function have over the set of complex numbers?

OpenStudy (aum):

Apply Descarte's Rule of Signs.

OpenStudy (samsan9):

is that the one with the sign changes?

OpenStudy (aum):

yes.

OpenStudy (anonymous):

@aum I feel like I did something wrong can you please check

OpenStudy (aum):

\(f(x) = (x^{3}-x^{2})-(4x+4)\) (4x+4) instead of (4x-4)

OpenStudy (samsan9):

what does find the zeroes mean? does that mean to plug in the zeros into the variables?

OpenStudy (aum):

zeros are the x values that will make the function zero. If you graph the function, it will cross the x axis at those zeros.

OpenStudy (anonymous):

sam can you check your equation again? : o make sure the signs are correct?

OpenStudy (samsan9):

ok hold up

OpenStudy (samsan9):

sorry it was supposed to be a +4

OpenStudy (anonymous):

oh then I was right! xD

OpenStudy (anonymous):

\[f(x)= (x^{3} - x^{2}) -(4x-4)\] \[f(x) = x^{2}(x-1) -4(x-1)\] \[f(x)= (x^{2}-4)(x-1)\] then find zeros

OpenStudy (aum):

If f(x) can be factored, no need for Descarte's Rule of Signs. Factor, equate it to zero and solve for x.

OpenStudy (samsan9):

so x^2-4=0 and x-1=0?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Thank you @aun, you're super smart! I'll look up that rule :)

OpenStudy (samsan9):

i got x=2 and x=1

OpenStudy (aum):

Thanks study100. You are smart for catching the sign error in the original posting.

OpenStudy (anonymous):

x^2 = 4 x= plus minus 2

OpenStudy (samsan9):

yeah this is the second time lol sorry about that and and oh yeah once its square rooted that means its like plus or minus everything right?

OpenStudy (anonymous):

Yes :) so 3 zeros

OpenStudy (samsan9):

yes thanks to both of you